OpenStudy (anonymous):

2x²+14x-60=0 First time using Open Study please help :)

OpenStudy (aravindg):

Welcome to Openstudy!! First you can divide the entire equation by 2.

OpenStudy (aravindg):

What do you get? :)

OpenStudy (anonymous):

8 I am new to math and really need this to impress my frinds

OpenStudy (aravindg):

Good. Can you divide the whole equation by 2?

OpenStudy (anonymous):

did n gut 8?

OpenStudy (aravindg):

:O How can you show?

OpenStudy (anonymous):

what?

OpenStudy (aravindg):

hw did u get 8?

OpenStudy (anonymous):

This is what you should have done. I am curious to know what you did to get 8? Hahaa well you first want to divide the whole equation by 2 like AracindG said so 2x^2 divided by 2=x^2 then 14x divided by 2=+7x and -60 divided by 2=-30 So the equation we now have is \[x^2+7x-30=0\]

OpenStudy (anonymous):

k

OpenStudy (aravindg):

Then you can use the quadratic formula.

OpenStudy (anonymous):

what that

OpenStudy (aravindg):

Do you know that?

OpenStudy (anonymous):

no

OpenStudy (aravindg):

http://www.purplemath.com/modules/quadform.htm

OpenStudy (anonymous):

cant click link from random man

OpenStudy (aravindg):

Why not?

OpenStudy (anonymous):

It's purplemath my math teacher used to send us there all the time. It's a math site that just helps you with equations and gives examples.

OpenStudy (anonymous):

\[x=\frac{ -b \pm \sqrt{b^2 -4ac} }{ 2a }\]

OpenStudy (anonymous):

A quadratic equation like the one you have looks like this \[ax^2+bx+c=0\]

OpenStudy (anonymous):

So in you case a=1 b=7 c=-30

OpenStudy (anonymous):

Never mind I found what I needed you can stop trying to help me